OpenStudy (anonymous):

Use natural logarithms to solve the equation. Round to the nearest thousandth.

OpenStudy (anonymous):

\[3e ^{2x}+2=28\]

OpenStudy (anonymous):

@satellite73 @nikato @agent0smith @timo86m can any of you please help me??

OpenStudy (sleepyjess):

@Mertsj

OpenStudy (mertsj):

1. Subtract 2 from both sides 2. Divide both sides by 3 3. Take the natural log of both sides.

OpenStudy (anonymous):

i know how to do the first two steps but its the last one where i get confused!

OpenStudy (mertsj):

Use the property that: \[\ln e ^{2x}=2x \ln e\]

OpenStudy (mertsj):

Remember that the ln e = 1

OpenStudy (mertsj):

Divide both sides by 2

OpenStudy (sleepyjess):

And this is why I called you @Mertsj ... you helped me so much with logs the other day :-)

OpenStudy (mertsj):

@emma97 Do you have it now?

OpenStudy (anonymous):

what i have so far is: e^2x=8.667

OpenStudy (mertsj):

\[\ln e ^{2x}=\ln 26\] \[2x \ln e=3.25097\]

OpenStudy (mertsj):

\[2x=3.25097\]

OpenStudy (anonymous):

i got 1.625485 but thats not any of my answer choices!

OpenStudy (mertsj):

Did you round to the nearest thousandth as the directions say?

OpenStudy (anonymous):

are you asking if i rounded the last response i just put?

OpenStudy (anonymous):

or the one before?

OpenStudy (mertsj):

2x=3.258096538021482 x=1.629048269010741=1.629

OpenStudy (mertsj):

After you solve the problem you must round your answer to the nearest thousandth.

OpenStudy (anonymous):

here ill show you my choices! because there isn't any choices that look like that lol

OpenStudy (mertsj):

Oh no.; Hold the phone. I forgot to divide by 3

OpenStudy (anonymous):

yeah lol i was trying to figure out what happened hahah

OpenStudy (mertsj):

\[3e ^{2x}+2=28\] \[3e ^{2x}=26\] \[e ^{2x}=\frac{26}{3}=8.6666666666....\] \[\ln e ^{2x}=\ln 8.6666666666666...\] \[2x=2.15948424934568\] \[x=1.07974212467284=1.080\]

OpenStudy (anonymous):

so then its option a?

OpenStudy (mertsj):

Perhaps you posted the directions incorrectly because the answer choices are given to the nearest ten thousandth...and not the nearest thousandth.

OpenStudy (anonymous):

maybe i did! I'm not sure but those are my answer choices lol!

OpenStudy (mertsj):

So anyway, it's the first one.

OpenStudy (anonymous):

thank you so so so much! your amazing!

OpenStudy (mertsj):

yw